QUESTION 1 15 points Save An Calc?late the force (in N) needed to bring a 1000 kg car to rest from a speed of 90.0 km/h in a distance of 115 m (a fairly typical distance for a non-panic stop).
Answer #1

Net force is given by

F = m*a

m = 1000 kg

a = acceleration

V^2 = U^2 + 2*a*d

d = 115 m

V = 0 m/sec

U = 90 km/hr = 25 m/sec

a = (0^2 - 25^2)/(2*115)

a = -2.717 m/sec^2

So,

F = 1000*2.717 = 2717 N